# ASTGI: Adaptive Spatio-Temporal Graph Interactions for Irregular Multivariate Time Series Forecasting

### v2

@@ 3.4.1 Message Function
${\bm{m}}_{j\to i}^{(l)}=\text{MLP}_{\text{msg}}({\bm{h}}_{j}^{(l)}\oplus({\bm{p}}_{i}-{\bm{p}}_{j}))$ (4)
@@ 3.4.2 Aggregation Function
Next, node $i$ aggregates all incoming messages from its causal neighborhood $\mathcal{N}(i)$ through a weighted sum. The aggregation weights $a_{ij}$ are calculated via the dynamic scoring mechanism in Section 3.3, reflecting the relative importance of each neighbor in the current interaction. This aggregation operation can adaptively focus on the most informative neighbors.
${\bm{m}}_{i}^{(l)}=\sum_{j\in\mathcal{N}(i)}a_{ij}\cdot{\bm{m}}_{j\to i}^{(l)}$ (5)
@@ 3.4.3 Update Function
${\bm{h}}_{i}^{(l+1)}=\text{LayerNorm}({\bm{h}}_{i}^{(l)}+\text{MLP}_{\text{update}}({\bm{m}}_{i}^{(l)}))$ (6)
@@ 3.5 Query Point-based Prediction
We unify the prediction task into this framework. A prediction request for a target time $t_{q}$ and target variable $c_{q}$ is treated as a query point, and its value is predicted by applying regression to the aggregated information from its historical neighborhood. The key difference between this prediction process and the feature propagation stage lies in the network modules used. Instead of reusing the scoring and message networks from the feature propagation layers, we design a separate set of scoring and fusion networks specifically for the regression task. This parameter separation allows the model to optimize independently for two functionally distinct sub-tasks—iterative feature updating during multi-layer propagation and direct numerical regression at the final prediction step—thereby improving prediction flexibility and accuracy.
The prediction process is as follows:
@@ 3.5.2 Query Relation Scoring and Weighted Fusion
$s_{qi}=\text{MLP}_{\text{query\_score}}\left(({\bm{p}}_{q}-{\bm{p}}_{i})\oplus{\bm{h}}_{i}^{(L)}\right)$ (7)
These scores are normalized through a Softmax function to obtain a set of interaction weights $a_{qi}$. Finally, these weights are used to perform a weighted fusion of neighbor information to generate a fusion vector ${\bm{h}}_{q}$. Before aggregation, we also use a value network $\text{MLP}_{\text{value}}$ to transform the neighbor features to extract the most valuable information segments for the prediction:
${\bm{h}}_{q}=\sum_{i\in\mathcal{N}(q)}a_{qi}\cdot\text{MLP}_{\text{value}}({\bm{h}}_{i}^{(L)})$ (8)
@@ 3.6 Training Objective
$\mathcal{L}=\frac{1}{|{\mathbb{S}}_{\text{query}}|}\sum_{(t_{j},x_{j},c_{j})\in{\mathbb{S}}_{\text{query}}}(\hat{x}_{j}-x_{j})^{2}$ (9)
@@ 4.1 Experimental Setup
Datasets and Baselines. We conduct experiments on four widely-used public IMTS datasets: MIMIC, PhysioNet, Human Activity, and USHCN. To ensure fairness and comparability of the results, all datasets are preprocessed by strictly following the standard procedures established in prior state-of-the-art works (Yalavarthi et al., 2024; Zhang et al., 2024). We uniformly split the data into training, validation, and test sets with a ratio of 80%, 10%, and 10%, respectively. We selected a total of twelve state-of-the-art models from two main categories designed for irregular time series as baselines for comparison. For a detailed description of the datasets and baselines, please refer to the Appendix B.1 and B.2.
Implementation Details. All our experiments were conducted on a server equipped with an NVIDIA A800 GPU and implemented using the PyTorch 2.6.0+cu124 framework. All models are trained using the Mean Squared Error (MSE) as the loss function and optimized with the AdamW optimizer. We set the maximum number of training epochs to 300 and employ an early stopping strategy, where training is terminated if the model’s performance on the validation set does not improve for 5 consecutive epochs. To ensure a fair comparison across all models, we primarily adopted the hyperparameter settings reported in the original papers for the baseline models. Building on these configurations, we conducted further search and fine-tuning of key hyperparameters on the validation set for some models to ensure that each achieved a competitive level of performance. To ensure reproducibility and mitigate the effects of randomness, each experiment is run independently with five different random seeds (from 2024 to 2028), and we report the mean and standard deviation. Detailed hyperparameter configurations for all models are provided in the Appendix B.2.
@@ 4.2 Main Results
4.3 Ablation Study
Dataset Human Activity USHCN PhysioNet MIMIC
Metric MSE MAE MSE MAE MSE MAE MSE MAE
@@ 4.3 Ablation Study
Table 2: Ablation study of ASTGI components. Results are reported in MSE and MAE (mean ± std). The performance of our full model is highlighted in bold.
To verify the effectiveness of each component in the ASTGI framework, we conduct a series of ablation studies—see Table 2. We draw the following four key conclusions. (1) the learnable coordinate space enhances representation power. Replacing the learnable time and channel embeddings with fixed, non-parametric encodings leads to a significant drop in performance. This indicates that an adaptively learned metric space is crucial for capturing the unique non-linear patterns and inter-variable correlations within the data. (2) the data-driven adaptive graph is superior to a static structure. Degrading the basis for neighborhood search from the learned spatio-temporal coordinates to the original timestamps results in a notable performance decline. This demonstrates that adaptively discovering neighbors in the learned metric space is more effective than relying on fixed rules of temporal proximity. (3) the relation-aware propagation mechanism improves interaction precision. Removing the spatio-temporal displacement vector (${\bm{p}}_{i}-{\bm{p}}_{j}$) as input when calculating interaction weights and messages causes a significant performance degradation. This highlights that modulating information based on the relative spatio-temporal position of neighbors is vital for capturing relation-dependent dynamics. (4) the dedicated query aggregation mechanism outperforms simple pooling. Replacing the interaction-weighted fusion in the prediction stage with simple neighborhood average pooling leads to a substantial drop in performance. This confirms that differentially weighting neighbor information based on its spatio-temporal relationship to the query point allows the model to focus more effectively on critical information during prediction.
(a) #Candidate Neighbors
(b) #Propagation Layers
(c) Channel embed_dim
(d) Hidden Dimension
4.4 Parameter Sensitivity

### v3

@@ 3.4.1 Message Function
${\bm{m}}_{j\to i}^{(l)}=\text{MLP}_{\text{msg}}({\bm{h}}_{j}^{(l)}\oplus({\bm{p}}_{i}-{\bm{p}}_{j})).$ (7)
@@ 3.4.2 Aggregation Function
Next, node $i$ aggregates all incoming messages from its causal neighborhood $\mathcal{N}(i)$ through a weighted sum. The aggregation weights $a_{ij}$ are calculated via the dynamic scoring mechanism in Section 3.3, reflecting the relative importance of each neighbor in the current interaction. This aggregation operation can adaptively focus on the most informative neighbors:
${\bm{m}}_{i}^{(l)}=\sum_{j\in\mathcal{N}(i)}a_{ij}\cdot{\bm{m}}_{j\to i}^{(l)}.$ (8)
@@ 3.4.3 Update Function
${\bm{h}}_{i}^{(l+1)}=\text{LayerNorm}({\bm{h}}_{i}^{(l)}+\text{MLP}_{\text{update}}({\bm{m}}_{i}^{(l)})).$ (9)
@@ 3.5 Query Point-based Prediction
We unify the prediction task into this framework. A prediction request for a target time $t_{q}$ and target variable $c_{q}$ is treated as a query point, and its value is predicted by applying regression to the aggregated information from its historical neighborhood. The key difference between this prediction process and the feature propagation stage lies in the network modules used. Instead of reusing the scoring and message networks from the feature propagation layers, we design a separate set of scoring and fusion networks specifically for the forecasting task. This parameter separation allows the model to optimize independently for two functionally distinct sub-tasks—iterative feature updating during multi-layer propagation and direct numerical regression at the final prediction step—thereby improving prediction flexibility and accuracy.
@@ 3.5.2 Query Relation Scoring and Weighted Fusion
$s_{qi}=\text{MLP}_{\text{query\_score}}\left(({\bm{p}}_{q}-{\bm{p}}_{i})\oplus{\bm{h}}_{i}^{(L)}\right).$ (10)
These scores $s_{qi}$ are normalized through a Softmax function to obtain a set of interaction weights $a_{qi}$. Finally, these weights are used to perform a weighted fusion of neighbor information to generate a fusion vector ${\bm{h}}_{q}$. Before aggregation, we also use a value network $\text{MLP}_{\text{value}}$ to transform the neighbor features to extract the most valuable information segments for the prediction:
${\bm{h}}_{q}=\sum_{i\in\mathcal{N}(q)}a_{qi}\cdot\text{MLP}_{\text{value}}({\bm{h}}_{i}^{(L)}).$ (11)
@@ 3.6 Training Objective
$\mathcal{L}=\frac{1}{|{\mathbb{S}}_{\text{query}}|}\sum_{(t_{j},x_{j},c_{j})\in{\mathbb{S}}_{\text{query}}}(\hat{x}_{j}-x_{j})^{2},$ (12)
@@ 4.1 Experimental Setup
Datasets and Baselines. We conduct experiments on four widely-used public IMTS datasets spanning diverse application domains: MIMIC, PhysioNet, Human Activity, and USHCN. A statistical overview of these datasets is provided in Table 1. These four benchmarks collectively cover a broad spectrum of real-world irregularity characteristics, ranging from high-dimensional clinical records (MIMIC with 96 variables) to low-dimensional climate monitoring (USHCN with 5 variables), and from densely sampled motion signals (Human Activity) to sparsely observed physiological indicators (MIMIC). To ensure fairness and comparability of the results, all datasets are preprocessed by strictly following the standard procedures established in prior state-of-the-art works (Yalavarthi et al., 2024; Zhang et al., 2024b). We uniformly split the data into training, validation, and test sets with a ratio of 80%, 10%, and 10%, respectively. We selected a total of twelve state-of-the-art models from two main categories designed for irregular time series as baselines for comparison. For a detailed description of the datasets and baselines, please refer to the Appendix B.1 and B.2.
Table 1: Statistical overview of the four benchmark datasets used in our experiments.
Property MIMIC PhysioNet Human Activity USHCN
Domain Healthcare Healthcare Biomechanics Climate
# Samples 21,250 11,981 1,359 1,114
# Variables 96 36 12 5
Max Sequence Length 96 47 131 337
Avg # Observations 144.6 308.6 362.2 313.5
Implementation Details. All our experiments were conducted on a server equipped with an NVIDIA A800 GPU and implemented using the PyTorch 2.6.0+cu124 framework. All models are trained using the Mean Squared Error (MSE) as the loss function and optimized with the AdamW optimizer. We set the maximum number of training epochs to 300 and employ an early stopping strategy, where training is terminated if the model’s performance on the validation set does not improve for 5 consecutive epochs. To ensure a fair comparison across all models, we primarily adopted the hyperparameter settings reported in the original papers for the baseline models. Building on these configurations, we conducted further search and fine-tuning of key hyperparameters on the validation set for some models to ensure that each achieved a competitive level of performance. To ensure reproducibility and mitigate the effects of randomness, each experiment is run independently with five different random seeds, and we report the mean and standard deviation. Detailed hyperparameter configurations for all models are provided in the Appendix B.2.
@@ 4.2 Main Results
(a) K
(b) L
(c) embedding d_c
(d) d_model
4.3 Parameter Sensitivity
4.4 Ablation Studies
Table 3: Ablation study of ASTGI components. Results are reported in MSE (mean $\pm$ std). The performance of our full model is highlighted in bold.
Model / Dataset Human Activity USHCN PhysioNet MIMIC
@@ 4.4 Ablation Studies
Table 4: Ablation study of ASTGI components. Results are reported in MAE (mean $\pm$ std). The performance of our full model is highlighted in bold.
Model / Dataset Human Activity USHCN PhysioNet MIMIC
[…]
To verify the effectiveness of each component in the ASTGI framework, we conduct a series of ablation studies—see Table 3 for the MSE metric and Table 4 for the MAE metric. We draw the following four key conclusions. (1) the learnable coordinate space enhances representation power. Replacing the learnable time and channel embeddings with fixed, non-parametric encodings leads to a significant drop in performance. This indicates that an adaptively learned metric space is crucial for capturing the unique non-linear patterns and inter-variable correlations within the data. (2) the data-driven adaptive graph is superior to a static structure. Degrading the basis for neighborhood search from the learned spatio-temporal coordinates to the original timestamps results in a notable performance decline. This demonstrates that adaptively discovering neighbors in the learned metric space is more effective than relying on fixed rules of temporal proximity. (3) the relation-aware propagation mechanism improves interaction precision. Removing the spatio-temporal displacement vector (${\bm{p}}_{i}-{\bm{p}}_{j}$) as input when calculating interaction weights and messages causes a significant performance degradation. This highlights that modulating information based on the relative spatio-temporal position of neighbors is vital for capturing relation-dependent dynamics. (4) the dedicated query aggregation mechanism outperforms simple pooling. Replacing the interaction-weighted fusion in the prediction stage with simple neighborhood average pooling leads to a substantial drop in performance. This confirms that differentially weighting neighbor information based on its spatio-temporal relationship to the query point allows the model to focus more effectively on critical information during prediction.
@@ 5 Conclusion
Acknowledgements
This work was partially supported by the National Natural Science Foundation of China (No. 62372179, No. 62472174) and the Fundamental Research Funds for the Central Universities. Jilin Hu is the corresponding author of the work.
@@ Acknowledgements
Reproducibility statement
The promise that all experimental results can be reproduced. We have released our model code in an anonymous repository: https://github.com/decisionintelligence/ASTGI.